HUBBLE SPACE TELESCOPE DAILY REPORT #2411 PERIOD COVERED: 0000Z (UTC) 07/09/99 - 0000Z (UTC) 07/12/99 Daily Status Report as of 193/0000Z 1.1 Completed Two Sets of STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.2 Completed Three Sets of WF/PC-2/STIS/CCD 8267 (Taking the Measure of Planets in the Globular Cluster 47 Tucanae) Both the WF/PC-2 and the Space Telescope Imaging Spectrograph (CCD) were used to make observations of the globular cluster NGC-104 (a.k.a., Globular Cluster 47 Tucanae) to search for stars with planetary companions in close orbits. HST is uniquely capable of providing the requisite, high-dynamic range, time-series photometry. These observations extended over a several day period. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed STIS/CCD 8412 (STIS CCD Imaging Flats C8) The Space Telescope Imaging Spectrograph (CCD) was used to investigate the flat-field stability of the unit over a monthly period. A bias calibration observation was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed STIS/CCD 8409 (STIS CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and to track the evolution of hot columns. The observations were executed as scheduled, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 54 Successful : 54 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Operations Notes: The nine-day “Tucanae” proposal completed this weekend. There were no problems encountered in either the acquisition support or in the extensive data collection activities. The SSR EDAC error counter was cleared thirteen times per ROP SR-1. Per an operations note, the K-based recharge ratio for Sun/orbit time capacity testing was adjusted beginning on day 191. The STIS MCE-1 reset at 191/010559Z while the low voltage was on and while in an SAA interval. The STIS flight software error counter was reset at 191/0126Z per ROP NS-12. The MAMA-1 was recovered at 193/0229Z via SMS nominal SMS commanding occurring at the time of a high voltage on event. The SSR pointers were set at 192/1632Z as directed by ROP SR-3. The NSSC-1 status buffer was dumped at 192/1650Z per ROP NS-3. As described in HSTAR 7190, the STIS EMC experienced a re-try at 193/060219Z, indicating the Echelle/Mode/Calibration (EMC) mechanism had initially failed to move. Again, the STIS flight software error counter was reset per ROP NS-12. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH